a man is 21 years older than his son. in six years, the man will be 5 times tha age of his son. what is age of the son now?
Answer by Alan3354(69443)

a man is 21 years older than his son. in six years, the man will be 5 times tha age of his son. what is age of the son now?
-----------
f = s + 21
f+6 = 5*(s+6)
-----
Sub for f, the father's age
s + 21 + 6 = 5s + 30
s + 27 = 5s + 30
4s = -3
s = -3/4 (son's age)
= -9 months
------
His son is not yet born, but might be a zygote.